OpenStudy (anonymous):

find the stationary point of the curve y=-xlogx and also determine the nature of the stationary point. please show all working out

OpenStudy (anonymous):

dy/dx = -x * 1/x + -1 logx (assuming its log to base e) so -1 - 1 logx = 0 at stationary points -1 = log x x = e^-1 stationary point at (e^-1, -0.3679) second derivative = -1 * 1/x which is negative stationary point is a maximum

OpenStudy (anonymous):

sorry the point is (0.3679,0.3679)

OpenStudy (anonymous):

or in terms of e it is (e^-1,e^-1)

OpenStudy (anonymous):

ok with that virtus?

OpenStudy (anonymous):

i used the product rule to differentiate

OpenStudy (anonymous):

@ jimmyrep thanks a bunch but could you just run through how to determine the nature of stationary points like the steps cause i am quite confused on that part thanks

OpenStudy (anonymous):

and can you also brief me on how to find points of inflexions thanks

OpenStudy (anonymous):

ok - if the second derivative is negative its a maximum positive its minimum here the second derivative = d(-1 - 1 logx) dx = -1/x plug in x = 0.3679 : it is -1/0.3679 which is negative so maximum the rate of change of dy/dx going over the 'hump' in a maximum is towards the negative gradient and with minimum its the opposite

OpenStudy (anonymous):

ok if its a point of inflection that is horizontal then second derivative will equal zero other points of inflection are more complicated - sorry i have no time to go into those - i have to go out.

OpenStudy (anonymous):

oh thank you jimmyrep for the much informative explanation! =D